# supplementary

### v1

@@ 5.1. Data Sources and Variables
We conduct a spatial and non-spatial analysis of US county-level associations between structural characteristics and three health outcomes: obesity prevalence, diabetes prevalence, and diabetes-related cancer mortality. Obesity and diabetes are both associated with increased risk of cancer, treatment complications, and higher mortality rate (Harborg et al., 2024). On the other hand, many epidemiological investigations often study relations between socioeconomic or environmental factors and a single health outcome such as obesity prevalence (see, e.g., Slack et al., 2014; Bennett et al., 2011; Salerno et al., 2024, and references therein), diabetes prevalence (e.g., Hill-Briggs et al., 2020; Cunningham et al., 2018), or cancer mortality rates (e.g., O’Connor et al., 2018; Dong et al., 2022). Furthermore, previous epidemiological studies have demonstrated the need to account for spatial autocorrelation on the county geographic level (e.g., Oh et al., 2024; Slack et al., 2014; Hipp and Chalise, 2015). We seek to consolidate the aforementioned analysis goals while accounting for spatial autocorrelation and more efficiently estimating associations according to the theoretical results of Section 3.2. Our secondary analysis goal is to assess differences in estimation and statistical inference between employing a non-spatial versus spatial analysis model.
Figure 2. US county-level obesity prevalence, diabetes prevalence, diabetes-related cancer mortality rates, and percent of people living in a food desert.
Estimates of diabetes and obesity age-adjusted prevalence rates among adults aged 20 years or more for 3,142 US counties in 2015 were extracted from the US Diabetes Surveillance System, a database maintained by the Centers for Disease Control and Prevention (2015) (CDC). We define diabetes-related cancer mortality with an underlying cause of death related to liver, pancreatic, endometrial, colorectal, breast, or bladder cancer (Giovannucci et al., 2010). We obtained age-adjusted mortality rates aggregated from 2010 to 2020 for 3,008 counties based on US resident death certificates classified with the corresponding ICD-10-CM (International Classification of Diseases, Tenth Revision, Clinical Modification) code from the 1999-2020 Underlying Cause of Death WONDER database, made publicly available by the Centers for Disease Control and Prevention, National Center for Health Statistics (2021).
Our analysis accounts for county-level characteristics across five established critical determinants of population health, which include the economic context, healthcare context, natural and built environment, educational context, and population structure (Slack et al., 2014; Myers et al., 2015, 2016). We account for the economic context of counties through the median household income, poverty rate, and unemployment rate in 2015. The healthcare context is represented by the percentage of the population under 65 years of age without health insurance, the number of primary care physicians per 100,000 people, and the number of outpatient visits per 100,000 people in 2015. Economic and healthcare variables were extracted from the US Department of Health and Human Services 2019-2020 County Level Area Health Resources Files (AHRF), which includes historical data from previous years.
The natural and built environment encapsulates both the food environment and the physical environment, otherwise known as the recreational context. We quantify the food environment by the percentage of the population living in a food desert and the percentage of the population receiving food stamp or Supplemental Nutrition Assistance Program (SNAP) benefits, both estimated for 2015. A person is defined as living in a food desert or low-access area if they live more than one mile from a supermarket or large grocery store in an urban area or more than ten miles in a rural area. Estimates of low access rates were extracted from the Food Environment Atlas, made publicly available by Economic Research Service , U.S. Department of Agriculture (USDA)(2020) (ERS) of the US Department of Agriculture (USDA). The percentages of food stamp or SNAP recipients were drawn from the 2019-2020 AHRF. The recreational context was captured by the age-adjusted percentage of adults aged 18 years or older in the US who are physically inactive in 2015, and the number of recreation and fitness facilities per 1,000 people in 2016. A person is defined as physically inactive if they obtain less than 10 minutes per week of moderate or vigorous activity in each category of physical activity of work, leisure time, and transportation. Data on physical inactivity were drawn from the CDC’s US Diabetes Surveillance System, while data on recreation and fitness facilities was taken from the USDA’s Food Environment Atlas.
The population structure was captured by six variables: the percentage of the population that was non-Hispanic black, the percentage of the population that was Hispanic, the percentage of the population at least 65 years old, the percentage of the population at most 18 years old, the percentage of the population living in urban areas, and the percentage of adults at least 25 years old that hold a high school diploma. Population structure data was extracted from the 2019-2020 AHRF, where all covariates except education were originally sourced from the 2010 Census of the US Census Bureau. The percentage of adults at least 25 years old that hold a high school diploma was sourced from the US Census Bureau’s 2011-2015 American Community Survey’s five-year estimates.
5.2. Statistical Analysis
@@
After aggregating all variables into one dataset, we drop counties with missing covariate data and subset to a contiguous region consisting of $n=2,960$ US counties for simplicity of analysis. For subsequent analysis, we center and scale the data such that each outcome and covariate has sample mean and standard deviation equal to zero and one, respectively. The three health outcomes as well as food desert exposure are mapped on the final study region in Figure 2, indicating the presence of considerable spatial clustering, such as the high obesity prevalence counties in the East North Central and East South Central states. We first conduct a non-spatial analysis using the non-spatial model in (4). Following the prior recommendations of Jin et al. (2007), we set $v=3$ and $\Sigma_{0}$ as a diagonal matrix where the $j$th diagonal element is $||(I_{n}-H)Y_{j}||^{2}/(n-p-1)$. We obtain 10,000 exact posterior samples of $\beta_{NS}$ and $\Sigma_{NS}$ by simulating $\Sigma_{NS}\sim\mbox{IW}(v_{\star},\Sigma_{\star})$ and then drawing one instance of $\beta_{NS}\sim\mbox{N}_{pk}\left(\operatorname{vec}\left(\widehat{\beta}_{NS}\right),\Sigma_{NS}\otimes(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}\right)$ for each sampled value of $\Sigma_{NS}$.
Turning to spatial analysis, we first assess spatial autocorrelation by computing Moran’s $I$ and Geary’s $C$ using the residuals of an ordinary least squares linear model for each outcome (Banerjee et al., 2015). Each statistic is significant under 10,000 random permutations (p-value = 1/10,001), indicating that the estimates from the non-spatial analysis model do not account for significant spatial dependence in all three health outcomes. To implement (6), we set the joint prior as in (7) with $v=3$, $\lambda_{R}=0.001$, and $\Sigma_{0}$ identical to the non-spatial model. We place a spatial CAR prior on the random effects by setting $W_{\phi}^{-1}=c(D_{W}-\alpha W)$ with the neighbor and adjacency matrices of the study region, $\alpha=0.99$, and $c=0.3963$. Although there are numerous choices for the spatial prior, we obtained a better fit quantified by a lower DIC using this CAR prior than a SAR spatial prior as in the simulation experiment with $W_{\phi}^{-1}=c(I_{n}-\alpha\widetilde{W})(I_{n}-\alpha\widetilde{W})^{\mathrm{\scriptscriptstyle T}}$, $\alpha=0.99$ and $c=41.41$.
@@
We employ the Metropolis within Gibbs sampling scheme detailed in Appendix D with proposal hyperparameters $s_{1}=0.05$, $s_{2}=0.05$, and $s_{3}=0.45$ to obtain $10,000\times 4\;\text{chains}=40,000$ posterior samples of $\{\beta_{S},\gamma,M,R\}$. One sample is saved every five update loops to reduce sampling autocorrelation and conserve computer memory. MCMC convergence is assessed via the $\widehat{R}$ and effective sample size statistics computed using the rstan package in R. The $\widehat{R}$ statistic is below the recommended threshold of 1.01 and the effective sample size exceeds 1000 for all sampled parameters (Vehtari et al., 2021). Coefficients are deemed statistically significant if the corresponding 95% highest posterior density credible interval does not contain zero.
5.3. Results
@@
Table 2 lists the standardized coefficient estimates for the spatial analysis model, indicating that diabetes prevalence in US counties has a moderate positive association with higher percentages of adults that are physically inactive ($\widehat{\beta}_{S}=0.47$; 95% CI: 0.44, 0.50) and a weak association with percent of non-Hispanic black residents ($\widehat{\beta}_{S}=0.22$; 95% CI: 0.17, 0.27). Other statistically significant predictors of diabetes prevalence include median income, unemployment rate, outpatient visit rate, percentage SNAP recipients, density of recreation facilities, percentage of residents at least 65 years of age, percentage of residents at most 18 years of age, and percentage of residents in urban areas. Outpatient visit rate and percent of residents 18 years old or younger are significant in the spatial model but not in the non-spatial model, whereas percentage of Hispanic residents is significant in the non-spatial model but not in the non-spatial model (Table 3).
Higher county obesity prevalence rates correlates with higher rates of physical inactivity ($\widehat{\beta}_{S}=0.47$; 95% CI: 0.44, 0.50), higher proportion of non-Hispanic black residents ($\widehat{\beta}_{S}=0.11$; 95% CI: 0.07, 0.16), lower median income ($\widehat{\beta}_{S}=-0.15$; 95% CI: -0.21, -0.097), and uninsured rate ($\widehat{\beta}_{S}=-0.18$; 95% CI: -0.23, -0.13). Other significant independent variables weakly associated with higher obesity prevalence include density of primary care physicians, percentage of Hispanic residents, percentage of residents at least 65 years of age, percentage of residents at most 18 years of age, and percentage of residents in urban areas. Primary care physician density, percent of residents at least 65 years old, and percent of residents in urban areas are significant in the spatial model but not in the non-spatial model, while outpatient visit rate and percent of residents living in food deserts are significant in the non-spatial model but not in the spatial model.
For diabetes-related cancer, the set of significant predictors of county-level mortality was identical between the spatial and non-spatial analysis models. Higher mortality rates are weakly associated with percentage of SNAP recipients ($\widehat{\beta}_{S}=0.32$; 95% CI: 0.25, 0.40), median income ($\widehat{\beta}_{S}=-0.18$; 95% CI: -0.26, -0.099), rates of high school diploma attainment ($\widehat{\beta}_{S}=-0.12$; 95% CI: -0.19, -0.052), percentage of non-Hispanic black residents ($\widehat{\beta}_{S}=0.15$; 95% CI: 0.084, 0.21), and percentage of Hispanic residents ($\widehat{\beta}_{S}=-0.15$; 95% CI: -0.22, -0.074). Other significant predictors weakly associated with diabetes-related cancer mortality include percent without health insurance, percentage of adults that are physically inactive, percentage of residents at least 65 years of age, percentage of residents at most 18 years of age, and percentage of residents in urban areas.
Variable Diabetes Obesity Cancer Mortality
Intercept 5.9e-06 (-0.022, 0.024) -1.9e-04 (-0.024, 0.025) -3.9e-05 (-3e-02, 3e-02)
Poverty Rate -0.043 (-0.1, 0.017) -6e-02 (-0.12, 7.8e-03) -0.042 (-0.12, 0.037)
Median Income -0.049 (-0.099, 3.3e-04) -0.19 (-0.24, -0.13) -9e-02 (-0.16, -0.026)
Unemployment 0.1 (0.068, 0.14) -0.014 (-0.049, 0.024) -0.11 (-0.15, -0.064)
Uninsured Rate 0.027 (-7.7e-03, 6e-02) -0.29 (-0.32, -0.25) -0.058 (-0.1, -0.014)
PCP Density -0.016 (-0.046, 0.015) -0.016 (-0.047, 0.017) -0.069 (-0.11, -0.028)
Outpatient Visits 0.023 (-3.2e-03, 0.049) 0.032 (4.2e-03, 6e-02) 0.042 (9.7e-03, 0.079)
Low Access 0.015 (-0.011, 0.038) 0.046 (0.021, 0.073) 0.047 (0.015, 8e-02)
SNAP Assistance 0.17 (0.12, 0.22) 2e-03 (-0.054, 0.053) 0.31 (0.25, 0.38)
Physical Inactivity 0.51 (0.48, 0.53) 0.6 (0.58, 0.63) 0.089 (0.054, 0.12)
Recreation Facilities -0.057 (-0.084, -0.029) -2e-02 (-0.049, 8.4e-03) -0.028 (-0.064, 7.5e-03)
Percent NH-Black 0.19 (0.16, 0.22) 0.061 (0.027, 0.092) 0.14 (0.1, 0.19)
Percent Hispanic -0.084 (-0.12, -0.051) -0.1 (-0.14, -0.068) -0.16 (-0.2, -0.11)
Percent $\geq$ 65 years -0.11 (-0.15, -0.074) 5.5e-03 (-0.032, 0.044) 0.099 (0.052, 0.15)
Percent $\leq$ 18 years 8.5e-03 (-0.023, 0.043) 0.16 (0.12, 0.19) 0.072 (0.032, 0.12)
Urban Percent 0.11 (0.073, 0.14) 0.035 (-2.1e-03, 0.071) 0.092 (0.047, 0.14)
HS Diploma Rate 8.1e-03 (-0.034, 0.053) 0.028 (-2e-02, 0.072) -0.21 (-0.27, -0.15)
Table 3. Non-spatial model regression coefficient posterior means and 95% highest posterior density credible regions.

### v2

@@ 5. Obesity, Diabetes, and Diabetes-Related Cancer Mortality in US Counties
We conduct a spatial and non-spatial analysis of US county-level associations between structural characteristics and three health outcomes: obesity prevalence, diabetes prevalence, and diabetes-related cancer mortality. Obesity and diabetes are associated with increased risk of cancer, treatment complications, and higher mortality rate (Harborg et al., 2024). However, many epidemiological investigations often study relations between socioeconomic or environmental factors and a single health outcome such as obesity prevalence (see, e.g., Slack et al., 2014; Bennett et al., 2011; Salerno et al., 2024, and references therein), diabetes prevalence (e.g., Hill-Briggs et al., 2020; Cunningham et al., 2018), or cancer mortality rates (e.g., O’Connor et al., 2018; Dong et al., 2022). Furthermore, previous epidemiological studies have demonstrated the need to account for spatial autocorrelation on the county geographic level (e.g., Oh et al., 2024; Slack et al., 2014; Hipp and Chalise, 2015). We seek to consolidate the aforementioned analysis goals while accounting for spatial autocorrelation and more efficiently estimating associations according to the results in Section 3.3. In Section LABEL:append:RDA_setup, we contrast our results to an analysis using the non-spatial model in (4).
Figure 2. Estimates of US county-level obesity prevalence in 2015, diabetes prevalence in 2015, diabetes-related cancer mortality rates from 2010 to 2020, and percentage of residents that were Hispanic in 2010. Counties with a missing covariate or response are shown in gray and excluded from the analysis.
@@
Our analysis accounts for county-level characteristics across five established critical determinants of population health, which include the economic context, healthcare context, natural and built environment, educational context, and population structure (Slack et al., 2014; Myers et al., 2015, 2016). For brevity, we refer to Section LABEL:append:RDA_setup for the original data sources and descriptions of the response and covariate data. After aggregating all variables into one dataset, we drop counties with missing covariate data and subset to a contiguous region consisting of $n=2,960$ US counties for simplicity of analysis. For subsequent analysis, we center and scale the data such that each outcome and covariate has sample mean and standard deviation equal to zero and one, respectively. The three health outcomes as well as county proportions of Hispanic residents are mapped on the final study region in Figure 2, indicating the presence of considerable spatial clustering, such as the high obesity prevalence counties in the East North Central and East South Central states.
@@
We first assess spatial autocorrelation by computing Moran’s $I$ and Geary’s $C$ using the residuals of an ordinary least squares linear model for each outcome (Banerjee et al., 2015). Each statistic is significant under 10,000 random permutations (p-value = 1/10,001), indicating significant spatial dependence in all three health outcomes. To implement (6), we set the joint prior as in (7) with $v=3$, $\lambda_{R}=0.01$, and $\Sigma_{0}$ as a diagonal matrix where the $j$th diagonal element is $(I_{n}-H)Y_{j}||^{2}/(n-p-1)$ following the prior recommendations of Jin et al. (2007). We place a spatial CAR prior on the random effects by setting $W_{\phi}^{-1}=c(D_{W}-\alpha W)$ with the neighbor and adjacency matrices of the study region, $\alpha=0.99$, and $c=0.3963$. This CAR prior results in a better fit quantified by a lower DIC compared to a SAR spatial prior. After standardizing all responses and predictors to have mean zero and unit variance, we employ the Metropolis within Gibbs sampling scheme detailed in Section LABEL:append:spatial_sampling_scheme with proposal hyperparameters $s_{1}=0.05$, $s_{2}=0.05$, and $s_{3}=0.45$ to obtain $10,000\times 4\;\text{chains}=40,000$ posterior samples of $\{B_{S},\gamma,M,R\}$. One sample is saved every five update loops to conserve computer memory. MCMC convergence is assessed via the $\widehat{R}$ and effective sample size statistics computed using the rstan package in R. The $\widehat{R}$ statistic is below the recommended threshold of 1.01 and the effective sample size exceeds 1000 for all sampled parameters (Vehtari et al., 2021). Coefficients are deemed statistically significant if the corresponding 95% highest posterior density credible interval does not contain zero.
@@
Table 4 lists the standardized coefficient estimates for the spatial analysis model, indicating that diabetes prevalence in US counties has a moderate positive association with higher percentages of adults that are physically inactive and a weak association with percent of non-Hispanic black residents. Other significant predictors of diabetes prevalence include median income, unemployment rate, outpatient visit rate, percentage SNAP recipients, density of recreation facilities, percentage of residents at least 65 years of age, percentage of residents at most 18 years of age, and percentage of residents in urban areas. County obesity prevalence rates correlates positively with higher rates of physical inactivity, higher proportion of non-Hispanic black residents, lower median income, and uninsured rate. Other significant independent variables weakly associated with higher obesity prevalence include density of primary care physicians, percentage of Hispanic residents, percentage of residents at least 65 years of age, percentage of residents at most 18 years of age, and percentage of residents in urban areas. Higher diabetes-related cancer mortality rates are weakly associated with percentage of SNAP recipients, median income, rates of high school diploma attainment, percentage of non-Hispanic black residents, percentage of Hispanic residents, percentage of residents without health insurance, percentage of adults that are physically inactive, percentage of residents at least 65 years of age, percentage of residents at most 18 years of age, and percentage of residents in urban areas.